OpenStudy (anonymous):

3-5.2p=-5.2p+3 is this conditional equation, a contradiction, or an identity. then describe the solution

OpenStudy (anonymous):

Are you able to solve for p on your own?

OpenStudy (anonymous):

Actually, just gather terms. Simplify the equation as much as you can.

OpenStudy (anonymous):

no these are hard to me

OpenStudy (anonymous):

Get all of the ps together.

OpenStudy (anonymous):

This is easy. It's an identity

OpenStudy (anonymous):

Relax, skippy. Let's help her understand why.

OpenStudy (anonymous):

lmao

OpenStudy (anonymous):

add 5.2p to both sides. 3 = 3

OpenStudy (anonymous):

I know you're excited that you know the answer.

OpenStudy (anonymous):

Smooth math, I post comments one at a time, let me do what I do, thanks

OpenStudy (anonymous):

I really want to understand how this process works I have a final in 5 weeks

OpenStudy (anonymous):

Roseanne, when you get something like this, begin by grouping like terms. Do you know what I mean by that?

OpenStudy (anonymous):

I apologize, Hero.

OpenStudy (anonymous):

Roseanne, I posted what to do above. Just simply add 5.2p to both sides and see what you get.

OpenStudy (anonymous):

Do you know what it means to group like terms?

OpenStudy (anonymous):

so combine terms means pull 5.2 and 5.2 together would i add them?

OpenStudy (anonymous):

yes hero I thought thank you

OpenStudy (anonymous):

very much

OpenStudy (anonymous):

add 5.2 to both sides of 5.2?

OpenStudy (anonymous):

gosh I want to understand this.

OpenStudy (anonymous):

Look at the equation 3-5.2p = -5.2- + 3. On the left side, there is one term with a p in it. On the right side there is one term with a p in it. You want to have all of the p terms on one side of the equation.

OpenStudy (anonymous):

I typed it wrong. It is 3-5.2p = -5.2p +3

OpenStudy (anonymous):

okay....

OpenStudy (anonymous):

So pick one of the p terms and find a way to move it over to the other side.

OpenStudy (anonymous):

Understand? What can you do?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

add 3 3 on left side and add 5.2 on left side to cancel?

OpenStudy (anonymous):

Yes, but since it's an equality, if you add something to one side, you have to add it to the other side too.

OpenStudy (anonymous):

im confused smooth

OpenStudy (anonymous):

So you have a p term on both sides, and you want p just on one side. So you decide to get rid of the p term on the left by adding 5.2p.

OpenStudy (anonymous):

You have to add 5.2p to both sides, not just 5.2 You don't add 5.2 separately

OpenStudy (anonymous):

3 - 5.2p = -5.2p +3 ^ ^ These are terms that are multiplied by the variable p. I want just one of them instead of 2.

OpenStudy (anonymous):

If I add 5.2p to the left side, it becomes 3 -5.2p + 5.2p = 3, which is what we want. But this is a balanced equation, so whatever we add to one side, we have to also add to the other.

OpenStudy (anonymous):

So we then add 5.2p to the right side also, and we get -5.2p +3 +5.2p = 3

OpenStudy (anonymous):

Which gives us the equation 3 = 3. Did you follow that?

OpenStudy (anonymous):

to be honest no these are so hard for me I dont know why

OpenStudy (anonymous):

Which part is confusing?

OpenStudy (anonymous):

maybe if we were one on one studying the comp thing is difficult

OpenStudy (anonymous):

Go to here. http://tinychat.com/mathisgreat

OpenStudy (anonymous):

3-5.2p=-5.2p+3 On the RHS move the +3 to the front of -5.2p. The result is: 3-5.2p= 3-5.2p Looks like an identity to me.